Given the Tc function Tc=10000 +100q-10q²+(1/3)q³ a. Find the MC function b. Find the slope of the MC function c. Find the output at which MC is equal to AVC
1
Expert's answer
2020-12-15T11:24:48-0500

a.


"MC=100-20q+q^2"

b.


"MC'=-20+2q"

c.


"AVC=100-10q+\frac{1}{3}q^2"


"100-20q+q^2=100-10q+\frac{1}{3}q^2"

"\frac{2}{3}q=10"


"q=15"